# ModernVBERT: Towards Smaller Visual Document Retrievers

### v1

@@ 3.2 Contrastive Training Design
Image resolution benefits are task-specific. Image resolution plays a critical role in VLM generative capabilities, notably in document-focused tasks, as higher-resolution inputs enables the model to capture finer visual cues (Hu et al., 2024a; Marafioti et al., 2025). Modality alignment is done at a fixed image resolution of 1024 x 1024 pixels and we report scores of contrastive training runs with varying settings in Table 1. Our findings confirm that embedding tasks are strongly sensitive to image-resolution. In particular, training with higher resolution inputs substantially improves the results on visual document retrieval benchmarks, consistent with prior work in generative settings Beyer et al. (2024); McKinzie et al. (2024). Furthermore, adding a cool-down phase by showing higher-resolution images towards the end of the modality alignment phase yields additional gains. This suggests that models can adapt their attention mechanisms to finer details when exposed to increased resolution. Interestingly, these findings do not hold in natural image tasks, where image up-scaling can even degrade performance.
@@
Increasing the pool of contrastive pairs. A severe limitation that current visual retrievers face is the lack of large volumes of high quality (document image, query pairs). Previous work (Ma et al., 2024; Faysse et al., 2025; Jiang et al., 2025; Zhang et al., 2025a) has relied on a mix of repurposed existing visual question answering datasets and synthetically generated queries with external LLMs. Even put together however, the field is only a year old, and these datasets remain small in size and often of poor quality.
A central question in our study is whether the abundance of text-only query–document pairs can be exploited to improve visual retrieval via cross-modal capability transfer. To probe this, we run contrastive training under three regimes. Unlike prior work that “warms up” visual retrievers or trains exclusively with text-only pairs (Ma et al., 2024; Jiang et al., 2024), we interleave text-only pairs and text–image pairs throughout training at a 1:1 ratio. The dataset sources are detailed in Appendix A.3.3
@@
Recipe. Putting together the results from our experiments, we devise a training recipe for a small visual document retriever ModernVBERT. It combines a state-of-the-art 150M text bidirectional encoder (Weller et al., 2025) with the ModernBERT architecture (Warner et al., 2024) and a small vision encoder SigLIP2-16B-512 of 100M parameters (Tschannen et al., 2025). We modality align both models with a MLM objective for 10B tokens, 3 times longer than during our experiments. To boost document understanding, we augment the input image resolution from 1024px to 2048px during a modality alignement cooldown stage (2B tokens). We call the resulting model ModernVBERT. Following the findings of Section 2, we then scale the contrastive training mix from previous experiments to combine document–query pairs with text-only pairs, and use 1 hard negatives for each document-query pair and 2 for each text-only pairs. We opt for a 2/1 text-to-image ratio following our ablation results introduced in Appendix C.2.1. This results in ColModernVBERT, a compact late interaction model. For reference, we also train BiModernVBERT, a single vector variant. More training details are provided in Appendix A.1.
@@
MoCa-3B (Chen et al., 2025) 3.75 80.1 53.8 66.9
GME-Qwen2 (Zhang et al., 2025a) 3.75 89.9 61.8 75.8
VLM2Vec (Jiang et al., 2025) 4.15 49.8 36.5 43.1
E5-V (Jiang et al., 2024) 8.36 62.7 49.4 56.1
ColPali (Faysse et al., 2025) 2.92 81.6 56.8 69.2 .222
ColQwen2.5 (Faysse et al., 2025) 3.75 89.5 61.5 75.5 .246
Jina-v4 (Günther et al., 2025) 3.75 90.4 60.1 75.2
NemoRetriever-3B (Xu et al., 2025) 4.40 91.0 66.3 78.7 .445
@@
Jina CLIP (Koukounas et al., 2024) 0.22 17.6 14.0 15.8 .023
BGE Visualized M3 (Zhou et al., 2024) 0.87 12.4 10.2 11.3 .018
SigLIP2-L-512/16 (Tschannen et al., 2025) 0.88 43.8 27.0 35.4 .004
ColFlor (Masry & Hoque, 2024) 0.17 68.8 43.0 55.9 .010
BiModernVBERT (ours) 0.25 63.6 35.7 49.7 .031
ColModernVBERT (ours) 0.25 81.2 56.0 68.6 .032
@@
Table 3: ViDoRe Leaderboard. Our model ColModernVBERT offers the best performance-size tradeoff, significantly outperforming existing sub-1B models and matching the performance of models up to 10x larger with substantially lower inference latency. Latency correspond to average query encoding latency on CPU. Models too large to run on standard CPUs are denoted with ✗. Bold values indicate the highest score among sub-1B models.
@@
ColModernVBERT. The resulting model, ColModernVBERT showcases strong performances on visual document retrieval benchmarks, especially relative to its size category (Figure 1). Despite having over 10 times less parameters than models such as ColPali released only a year ago, it is only 0.6 nDCG@5 points below on the aggregated ViDoRe benchmark scores (Table 3). It also edges many larger single-vector repurposed VLM models released within the year (Chen et al., 2025; Jiang et al., 2024; 2025). It however falls short of top model performance on ViDoRe which are built on larger decoder VLMs pretrained and aligned on billions of tokens of text and image data.
@@
Most sub-1B parameter models evaluated on document retrieval benchmarks are dual encoder models, since early fusion generative models that perform well are not common at this scale. The most related model is a 176M late interaction model, ColFlor (Masry & Hoque, 2024), trained from the Florence2 model (Xiao et al., 2023). ColFlor is 12.7 nDCG@5 points under ColModernVBERT. ColModernVBERT also largely outperforms off-the-shelf dual encoders, even when those have substantially larger parameter counts. These results highlights the benefits of multi-phase training and early fusion architectures for multi-modal document related tasks, even at smaller parameter counts. We also attribute the strong performance of ColModernVBERT at smaller model sizes to the symbiosis of native bidirectional attention and Late Interaction matching, which largely boosts performance relative to comparable decoder models (Section 3.1).
@@
Speed. As noted by Xiao et al. (2025b), multi-vector visual retrievers are not bottlenecked in their inference speed by the late interaction matching operation, but rather by the latency required to encode queries with the text model. Our model demonstrates that strong performance is not incompatible with speed, even when running inference on consumer CPUs, a standard setting in practical text retrieval applications. Latencies are computed by averaging query encoding times over 1000 varied queries, which are 60 character long on average, and ran on standard Google Colab CPU environments (Table 3). ModernVBERT achieves almost a 7x speedup on CPU over models with similar performances on ViDoRe, while models in its size category fall far behind.
@@ 5 Related Work
Repurposing VLMs for Representation Learning. Motivated by the zero-shot performances of generative VLMs (Alayrac et al., 2022; Lucas Beyer* et al., 2024; Bai et al., 2023), recent studies have explored repurposing these for multimodal embedding tasks (Ma et al., 2024; Faysse et al., 2025; Jiang et al., 2025; Zhang et al., 2025a). As backbone generative models improved, retriever performance improved as well showcasing the central impact of language model pretraining and modality alignement (Xu et al., 2025; Nussbaum et al., 2025). These model remain inherently constrained by their causal attention mechanisms which has been shown in text settings to limits represational efficiency  (Gisserot-Boukhlef et al., 2025; Weller et al., 2025). Recent work attempts to address this issue by modifying VLM attention during continual pretraining (Chen et al., 2025) or contrastive tuning  (Jiang et al., 2025; Xu et al., 2025), but no recent work attempts to align natively bidirectional language encoder models with vision encoders. The recent release of long sequence text encoders (Warner et al., 2024; Boizard et al., 2025) makes this possible.
Late Interaction in Visual Document Retrieval To further boost performance, visual document retrievers leverage the late interaction mechanism (Khattab & Zaharia, 2020) which matches multiple query embeddings with multiple document embeddings through the MaxSim operation  (Faysse et al., 2025; Günther et al., 2025; Xu et al., 2025). This enables more granular interactions between image and query tokens, at the cost of additional storage and a slight compute overhead during the matching operation. Efficiency gains have come from improving the storage costs through quantization (Bergum, 2025), token pruning (Faysse et al., 2024) and more recently the use of Matrioshka losses to compact multi-token representations (Xiao et al., 2025b). Ultimately, the performance bottleneck when running visual retrieval inference with such models now resides mostly in the necessity to rely on costly GPU hardware to encode queries, which sets apart text from vision retrieval. This paper fills this gap by using encoders that run on CPU, of parameter sizes comparable to commonly used local text embedding models (Chen et al., 2024; Enevoldsen et al., 2025).
@@ 6 Conclusion
Future Work & Limitations. By design, our analysis targets relatively small models. An important next step is to test whether the observed patterns persist at larger scales—for example, to more rigorously probe the interplay between late interaction and bidirectional attention. Our study also focuses exclusively on English. While we expect the broad trends to generalize and see clear value in releasing multilingual variants, it remains unclear how allocating parameters to additional languages trades off against the understanding of the vision modality, and to what extent this penalizes English retrieval performance as the number of languages are scaled (Fernandes et al., 2023). Finally, although we center on retrieval and sequence-level zero-shot classification, the modality-aligned encoder can be fine-tuned for a range of token-level tasks, including OCR error detection, token-level classification, visual named entity recognition, visually grounded token-level object detection, contextual embeddings (Conti et al., 2025). We release our base model to encourage exploration of these directions.
@@
Environmental Costs. Training ColModernVBERTrequired approximately 2,000 H100 GPU-hours in total, which we estimate corresponds to 41 kg of CO2555Carbon footprint estimated with Green Algorithms (Lannelongue et al., 2021): $E=t\times P\times\mathrm{PUE},\;\mathrm{CO_{2}e}=E\times\mathrm{CI}$. With $t=2000$ GPUh, $P=0.35$ kW (H100 PCIe), $\mathrm{PUE}=1.3$, and $\mathrm{CI}=45$ gCO2/kWh, this gives $E\approx 910$ kWh and $\mathrm{CO_{2}e}\approx 41$ kg., based on standard assumptions of GPU power draw, datacenter efficiency, and grid carbon intensity. This estimate follows methodologies such as Green Algorithms (Lannelongue et al., 2021) and related analyses of the carbon footprint of machine learning (Strubell et al., 2019; Patterson et al., 2021). Across the entire project, all combined experiments totaled about 18k H100-hours. To mitigate costs and promote sustainable research practices, we release all model checkpoints and training artifacts to facilitate reuse, extension, and reproducibility without necessitating retraining. Additionally, this work shows efficiency gains with smaller models to aim to limit the inference costs of visual retrieval, and consequently reduce the environmental footprint. Our model performs query encoding efficiently on CPUs, keeping inference costs low and reducing barriers to adoption.
@@
Detailed Contributions
PT is the first author of the project. Notably, he designed the modality alignement codebase, ran and supervised most large scale experiments across modality alignment and contrastive training, coordinated work, and was key in paper writing. QM ran large scale ablations on the data mixtures, contrastive training, and was responsible for the final training run. MC ran multiple experiments, including investigations into model merging and contributed to paper writing. AL focused on optimizing the data mixture and the contrastive training codebase. He was notably responsible for the initial cross-modality positive transfer results. GV and PC are senior contributors who helped with project framing, grant obtention, industry expertise and paper review. MF is last author and scientific lead of this work. He initiated and closely supervised the project from beginning to end, wrote the initial version of the multimodal contrastive training framework and greatly contributed to paper writing.
Acknowledgments
This work was carried out within the framework of the LIAGORA ”LabCom”, a joint laboratory supported by the French National Research Agency (ANR) and established between ILLUIN Technology and the MICS laboratory of CentraleSupélec. This work was performed using HPC resources from IDRIS with grant AD011016393. We warmly thank Hippolyte Gisserot-Boukhlef and Nicolas Boizard for sharing the controlled experiments LM checkpoints, Antoine Chaffin for his feedback on the modality alignment codebase and insights on Ettin’s modeling, as well as Andi Marafioti, Orr Zohar, and Miquel Farré for their valuable input and help on gathering the modality alignment dataset.
@@ A.1 Implementation and Resources
We list hyperparameters and resource details in Table 4 for the various training stages of our final models. We employ ZeRO stage 1 optimizer (Rajbhandari et al., 2020) for our modality alignment runs. All ablation models are contrastively trained with gradient checkpointing (Chen et al., 2016) to reduce memory usage. All training runs are performed with FlashAttention 2.0 (Dao, 2023). For LoRA configurations, we consistently use a rank r of $32$, lora_alpha of $32$, and a dropout of $0.1$. For the implementation, we start from m4666SmolVLM trainer, https://github.com/huggingface/smollm and ColPali [^7] https://github.com/illuin-tech/colpali codebases for training, and use the MTEB [^8] https://github.com/embeddings-benchmark/mteb repository for evaluation. [^9] We will release our training codebases in the public version of this paper

### v2

@@ 3.2 Contrastive Training Design
Image resolution benefits are task-specific. Image resolution plays a critical role in VLM generative capabilities, notably in document-focused tasks, as higher-resolution inputs enables the model to capture finer visual cues (hu2024mplugdocowl2highresolutioncompressingocrfree; marafioti2025smolvlmredefiningsmallefficient). Modality alignment is done at a fixed image resolution of 1024 pixels (longer side) and we report scores of contrastive training runs with varying settings in Table 1. To vary the resolution, images of the highest quality available are scaled to the desired size (often downscaled) before being fed to the image tokenizer. Our findings confirm that embedding tasks are strongly sensitive to image-resolution. In particular, training with higher resolution inputs substantially improves the results on visual document retrieval benchmarks, consistent with prior work in generative settings beyer2024paligemmaversatile3bvlm; mckinzie2024mm1methodsanalysis. Furthermore, adding a cool-down phase by showing higher-resolution images towards the end of the modality alignment phase yields additional gains. This suggests that models can adapt their attention mechanisms to finer details when exposed to increased resolution. Interestingly, these findings do not hold in natural image tasks, where high resolution can even degrade performance.
@@
Increasing the pool of contrastive pairs. A severe limitation that current visual retrievers face is the lack of large volumes of high quality (document image, query pairs). Previous work (ma2024unifyingmultimodalretrievaldocument; faysse2025colpaliefficientdocumentretrieval; jiang2025vlm2vectrainingvisionlanguagemodels; zhang2025gmeimprovinguniversalmultimodal) has relied on a mix of repurposed existing visual question answering datasets and synthetically generated queries with external LLMs. Even put together however, the field is only a year old, and these datasets remain small in size and often of poor quality.
A central question in our study is whether the abundance of text-only query–document pairs can be exploited to improve visual retrieval via cross-modal capability transfer. To probe this, we run contrastive training under three regimes. Unlike prior work that “warms up” visual retrievers or trains exclusively with text-only pairs (ma2024unifyingmultimodalretrievaldocument; jiang2024e5vuniversalembeddingsmultimodal), we interleave text-only pairs and text–image pairs throughout training at a 1:1 ratio. The dataset sources are detailed in Appendix A.3.3
@@
Recipe. Putting together the results from our experiments, we devise a training recipe for a small visual document retriever ModernVBERT. It combines a state-of-the-art 150M text bidirectional encoder (weller2025seqvsseqopen) with the ModernBERT architecture (modernbert) and a small vision encoder SigLIP2-16B-512 of 100M parameters (tschannen2025siglip2multilingualvisionlanguage). We modality align both models with a MLM objective for 10B tokens, 3 times longer than during our experiments. To boost document understanding, we augment the input image resolution from 1024px to 2048px during a modality alignment cooldown stage (2B tokens). We call the resulting model ModernVBERT. Following the findings of Section 2, we then scale the contrastive training mix from previous experiments to combine document–query pairs with text-only pairs, and use 1 hard negatives for each document-query pair and 2 for each text-only pairs. We opt for a 2/1 text-to-image ratio following our ablation results introduced in Appendix C.3.1. This results in ColModernVBERT, a compact late interaction model. For reference, we also train BiModernVBERT, a single vector variant. More training details are provided in Appendix A.1.
@@
MoCa-3B (chen2025mocamodalityawarecontinualpretraining) 3.75 80.1 53.8 66.9 158
VLM2Vec (jiang2025vlm2vectrainingvisionlanguagemodels) 4.15 49.8 36.5 43.1 211
GME-Qwen2 (zhang2025gmeimprovinguniversalmultimodal) 8.29 89.9 61.8 75.8 412
E5-V (jiang2024e5vuniversalembeddingsmultimodal) 8.36 62.7 49.4 56.1 434
ColPali (faysse2025colpaliefficientdocumentretrieval) 2.92 81.6 56.8 69.2 175
ColQwen2.5 (faysse2025colpaliefficientdocumentretrieval) 3.75 89.5 61.5 75.5 158
Jina-v4 (günther2025jinaembeddingsv4universalembeddingsmultimodal) 3.75 90.4 60.1 75.2 158
NemoRetriever-3B (xu2025llamanemoretrievercolembedtopperforming) 4.40 91.0 66.3 78.7 155
@@
Jina CLIP (koukounas_jina_2024) 0.22 17.6 14.0 15.8 14
BGE Visualized M3 (zhou2024vistavisualizedtextembedding) 0.87 12.4 10.2 11.3 38
SigLIP2-L-512/16 (tschannen2025siglip2multilingualvisionlanguage) 0.88 43.8 27.0 35.4 25
ColFlor (masrycolflor) 0.17 68.8 43.0 55.9 17
BiModernVBERT (ours) 0.25 63.6 35.7 49.7 20
ColModernVBERT (ours) 0.25 81.2 56.0 68.6 20
@@
Table 3: Performance on ViDoRe. Our model ColModernVBERT offers the best performance-size tradeoff, significantly outperforming existing sub-1B models and matching the performance of models up to 10x larger with substantially lower inference CPU latency Details and GPU latencies in Appendix C.6.2. Models marked with are not specifically trained for VDR. Bold values indicate the best performance amongst sub-1B models.
@@
ColModernVBERT. The resulting model, ColModernVBERT showcases strong performances on visual document retrieval benchmarks, especially relative to its size category (Figure 1). Despite having over 10 times less parameters than models such as ColPali released only a year ago, it is only 0.6 nDCG@5 points below on the aggregated ViDoRe benchmark scores (Table 3). It also edges many larger single-vector repurposed VLM models released within the year (chen2025mocamodalityawarecontinualpretraining; jiang2024e5vuniversalembeddingsmultimodal; jiang2025vlm2vectrainingvisionlanguagemodels). It however falls short of top model performance on ViDoRe which are built on larger decoder VLMs pretrained and aligned on billions of tokens of text and image data.
@@
Most sub-1B parameter models evaluated on document retrieval benchmarks are dual encoder models, since early fusion generative models that perform well are not common at this scale. The most related model is a 176M late interaction model, ColFlor (masrycolflor), trained from the Florence2 model (xiao2023florence2advancingunifiedrepresentation). ColFlor is 12.7 nDCG@5 points under ColModernVBERT. ColModernVBERT also largely outperforms off-the-shelf dual encoders, even when those have substantially larger parameter counts. These results highlights the benefits of multi-phase training and early fusion architectures for multi-modal document related tasks, even at smaller parameter counts. We also attribute the strong performance of ColModernVBERT at smaller model sizes to the symbiosis of native bidirectional attention and Late Interaction matching, which largely boosts performance relative to comparable decoder models (Section 3.1).
@@
Speed. As noted by xiao2025metaembedscalingmultimodalretrieval, multi-vector visual retrievers are not bottlenecked in their inference speed by the late interaction matching operation, but rather by the latency required to encode queries with the text model. Our model demonstrates that strong performance is not incompatible with speed, even when running inference on consumer CPUs, which is the standard setting in most industrial local deployments of text embedding models. Latencies are computed by averaging query encoding times of all NanoBEIR queries, which are 23.4 word and 147.5 character long on average, and are run with batch size 1 to replicate online use cases. To prevent RAM bottlenecks, we benchmark on very high RAM (2TB) CPU cloud environments, but note models larger than 3B parameter require more than 12 GB RAM to run optimally. [^7] With more standard CPU RAM settings such as those found in low-end servers or Google Colab (12GB RAM), models above 3B parameters must rely on memory offloading to run, which adds up to dozens of seconds of latency per query. (Table 3). ModernVBERT achieves more than a 7x speedup on CPU over models with similar performances on ViDoRe. We further report model latency results on GPU hardware in Appendix C.6.2. We notably demonstrate that with batched inference, ModernVBERT based query encoders are able to encode 5000 queries per second on Nvidia H100 GPUs. ModernVBert’s small model size also enables efficient batching when encoding documents.
@@ 5 Related Work
Repurposing VLMs for Representation Learning. Motivated by the zero-shot performances of generative VLMs (alayrac2022flamingovisuallanguagemodel; lucas_beyer_paligemma_2024; bai_qwen-vl_2023), recent studies have explored repurposing these for multimodal embedding tasks (ma2024unifyingmultimodalretrievaldocument; faysse2025colpaliefficientdocumentretrieval; jiang2025vlm2vectrainingvisionlanguagemodels; zhang2025gmeimprovinguniversalmultimodal). As backbone generative models improved, retriever performance improved as well showcasing the central impact of language model pretraining and modality alignment (xu2025llamanemoretrievercolembedtopperforming; nussbaum2025nomicembedtrainingreproducible). These model remain inherently constrained by their causal attention mechanisms which has been shown in text settings to limits represational efficiency  (gisserotboukhlef2025pretrainencodersmaskedlanguage; weller2025seqvsseqopen). Recent work attempts to address this issue by modifying VLM attention during continual pretraining (chen2025mocamodalityawarecontinualpretraining) or contrastive tuning  (jiang2025vlm2vectrainingvisionlanguagemodels; xu2025llamanemoretrievercolembedtopperforming), but no recent work attempts to align natively bidirectional language encoder models with vision encoders. The recent release of long sequence text encoders (modernbert; boizard2025eurobertscalingmultilingualencoders) makes this possible.
Late Interaction in Visual Document Retrieval To further boost performance, visual document retrievers leverage the late interaction mechanism (khattab_colbert_2020) which matches multiple query embeddings with multiple document embeddings through the MaxSim operation  (faysse2025colpaliefficientdocumentretrieval; günther2025jinaembeddingsv4universalembeddingsmultimodal; xu2025llamanemoretrievercolembedtopperforming). This enables more granular interactions between image and query tokens, at the cost of additional storage and a slight compute overhead during the matching operation. Efficiency gains have come from improving the storage costs through quantization (vespaScalingColPali), token pruning (faysse_croissantllm_2024) and more recently the use of Matrioshka losses to compact multi-token representations (xiao2025metaembedscalingmultimodalretrieval). Ultimately, the performance bottleneck when running visual retrieval inference with such models now resides mostly in the necessity to rely on costly GPU hardware to encode queries, which sets apart text from vision retrieval. This paper fills this gap by using encoders that run on CPU, of parameter sizes comparable to commonly used local text embedding models (chen_bge_2024; enevoldsen2025mmtebmassivemultilingualtext).
@@ 6 Conclusion
Future Work & Limitations. By design, our analysis targets relatively small models. An important next step is to test whether the observed patterns persist at larger scales—for example, to more rigorously probe the interplay between late interaction and bidirectional attention. Our study also focuses exclusively on English. While we expect the broad trends to generalize and see clear value in releasing multilingual variants, it remains unclear how allocating parameters to additional languages trades off against the understanding of the vision modality, and to what extent this penalizes English retrieval performance as the number of languages are scaled (pmlr-v202-fernandes23a). Finally, although we center on retrieval and sequence-level zero-shot classification, the modality-aligned encoder can be fine-tuned for a range of token-level tasks, including OCR error detection, token-level classification, visual named entity recognition, visually grounded token-level object detection, contextual embeddings (conti2025context). We release our base model to encourage exploration of these directions.
@@
Environmental Costs. Training ColModernVBERTrequired approximately 2,000 H100 GPU-hours in total, which we estimate corresponds to 41 kg of CO2888Carbon footprint estimated with Green Algorithms (Lannelongue2021GreenAlgorithms): $E=t\times P\times\mathrm{PUE},\;\mathrm{CO_{2}e}=E\times\mathrm{CI}$. With $t=2000$ GPUh, $P=0.35$ kW (H100 PCIe), $\mathrm{PUE}=1.3$, and $\mathrm{CI}=45$ gCO2/kWh, this gives $E\approx 910$ kWh and $\mathrm{CO_{2}e}\approx 41$ kg., based on standard assumptions of GPU power draw, datacenter efficiency, and grid carbon intensity. This estimate follows methodologies such as Green Algorithms (Lannelongue2021GreenAlgorithms) and related analyses of the carbon footprint of machine learning (Strubell2019Energy; Patterson2021Carbon). Across the entire project, all combined experiments totaled about 18k H100-hours. To mitigate costs and promote sustainable research practices, we release all model checkpoints and training artifacts to facilitate reuse, extension, and reproducibility without necessitating retraining. Additionally, this work shows efficiency gains with smaller models to aim to limit the inference costs of visual retrieval, and consequently reduce the environmental footprint. Our model performs query encoding efficiently on CPUs, keeping inference costs low and reducing barriers to adoption.
@@ A.1 Implementation and Resources
We list hyperparameters and resource details in Table 4 for the various training stages of our final models. We employ ZeRO stage 1 optimizer (rajbhandari2020zeromemoryoptimizationstraining) for our modality alignment runs. All ablation models are contrastively trained with gradient checkpointing (chen2016trainingdeepnetssublinear) to reduce memory usage. All training runs are performed with FlashAttention 2.0 (dao2023flashattention2fasterattentionbetter). For LoRA configurations, we consistently use a rank r of $32$, lora_alpha of $32$, and a dropout of $0.1$. For the implementation, we start from m4999SmolVLM trainer, https://github.com/huggingface/smollm and ColPali [^10] https://github.com/illuin-tech/colpali codebases for training, and use the MTEB [^11] https://github.com/embeddings-benchmark/mteb repository for evaluation. [^12] https://github.com/illuin-tech/modernvbert